# Markov Decision Processes                                  Under External Temporal Processes

### v2

@@ 4.2. Examples
Non-Markov self-exciting event processes that can be described as above are studied extensively in finance, economics, epidemiology, etc., in the form of variants of Hawkes and jump processes. While they are studied predominantly in continuous time and sometimes in discrete time, we consider discrete time versions to fit into the traditional discrete-time reinforcement learning framework.
Consider a discrete-time marked Hawkes process $\left(E_{t},X_{t}\right)_{n\in\mathbb{N}}$ defined as follows:
$\left(E_{t}\right)_{t\in\mathbb{N}}$ is a sequence of random variables taking values in $\left\{0,1\right\}$, with $E_{t}$ sampled from $\textsf{Bernoulli}\left(p_{t}\right)$, where the intensity $p_{t}$ at time $t$ depends on the realizations of $E_{t}$ at the previous time steps as:
@@
where $(\alpha_{t})$ and $(\beta_{t})$ are sequences that converge sufficiently quickly to zero.
The sum $S_{t}=\sum_{t^{\prime}=1}^{t}E_{t}$ is a self-exciting counting process that is the discrete-time counterpart of the continuous-time Hawkes process.
The above defined sequence of random variables satisfies Assumptions (i) and (ii) required for the external process in Section 4.1.1. $X_{t}$ is zero when no event occurs, and old events have a reduced influence on current events. More precisely, given two histories $(X_{t^{\prime}})_{t^{\prime}<t}$ and $(X^{\prime}_{t^{\prime}})_{t^{\prime}<t}$ that only differ by one event at some time $t^{\prime}\leq t-T$ with $E^{\prime}_{t^{\prime}}=X^{\prime}_{t^{\prime}}=0$ and $E_{t^{\prime}}=1$, $X_{t^{\prime}}=x$, the difference in intensities is $p_{t}-p^{\prime}_{t}=\alpha_{t-t^{\prime}}$, and so,
@@
For sufficiently quickly converging $\alpha_{T}$ and $\beta_{T}$, the total variation perturbation $N_{T}$ due to events older than $T$ time steps goes to zero, satisfying Assumption (ii).
While $(\alpha_{n})$ and $(\beta_{n})$ are any general sequences that need to satisfy some regularity conditions (Seol, 2015) and the above bound, they could also decay in a more structured manner. For instance, for some $c>0.5,\lambda>0$ letting
@@
5. Characterization of the problem setting
Given the MDP $(S,A,r,\gamma,Q)$ and an external discrete-time temporal process $\{(t_{i},X_{i})\}_{i}$ that acts on this MDP, we construct a new augmented MDP where the marks of all the events that occurred till time $t$ are appended to the state $s\in S$ to form a new augmented state $\bar{s}\in\bar{S}$, where
$\bar{s}=(s,x_{t},x_{t-1},x_{t-2},x_{t-3},\dots),\qquad\bar{S}=S\times\prod_{t=0}^{\infty}\mathcal{X}\left(\subset\mathbb{R}^{m}\times\prod_{t=0}^{\infty}\mathbb{R}^{d}\right).$
@@ 5. Characterization of the problem setting
Now, the state space of this MDP is infinite-dimensional and poses several challenges. However, since we assume that very old events do not affect the transition probabilities during the current time step, we should be able to find a suitable policy that only depends on a finite horizon of past events. This is formalized by the following result:
@@
Suppose the MDP and the external process satisfy the assumptions in 4.1.1 and 4.1.2.
@@
For each time $T>0$, we construct MDP $\mathcal{M}^{(T)}$ as having the same underlying stochastic process as $\mathcal{M}_{X}$, with only one difference: only events in the past $T$ time steps affect the current state transition, and all the events occurring before that have no effect and can effectively be replaced by zeroes. Depending on the context, the state space of this MDP can either be taken as $S\times\prod_{t=0}^{\infty}\mathcal{X}$ or $S\times\prod_{t=0}^{T}\mathcal{X}$, since all the older events are zeroed out.
Any policy $\pi$ for this MDP has an equivalent policy in $\mathcal{M}_{X}$ and vice-versa, where any extra elements in the state are ignored or considered zero.
@@
Let $\pi$ be an arbitrary policy as a function of the augmented state $\bar{s}\in\bar{S}$. That is, $\pi$ can be either stochastic or deterministic, and can depend on the current state $s\in S$ and any number of past events. Then,
$\left|\left|V^{\pi}_{\mathcal{M}^{(T)}}-V^{\pi}_{\mathcal{M}_{X}}\right|\right|_{\infty}\leq\frac{1}{1-\gamma}\left(||r||_{\infty}+\gamma||V^{\pi}_{\mathcal{M}^{(T)}}||_{\infty}\right)\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right),$
where $V^{\pi}_{\mathcal{M}}$ denotes the value function of policy $\pi$ in MDP $\mathcal{M}$.
@@
Part (1) follows from the assumptions and Theorem 3.1. For part (2), let $\pi^{*}$ be the deterministic optimal stationary policy for $\mathcal{M}_{X}$ and let $\pi^{*(T)}$ be the deterministic optimal stationary policy for $\mathcal{M}^{(T)}$. From Lemma 5.2, since $||r||_{\infty}\leq 1$ and $||V||_{\infty}\leq\frac{1}{1-\gamma}$,
$V^{\pi^{*(T)}}$ $\geq V^{\pi^{*(T)}}_{\mathcal{M}^{(T)}}-\frac{1}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$\geq V^{\pi^{*}}_{\mathcal{M}^{(T)}}-\frac{1}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
@@
Choosing such a $T$ gives the Theorem.
@@ 6. A policy iteration Algorithm
Now that an $\epsilon$-optimal policy exists that is a function of only a finite event horizon, we wish to find such a policy. We propose a policy iteration algorithm, which alternates between approximate policy evaluation and approximate policy improvement. The overall algorithm is given in 1.
@@
For the policy evaluation step, we only wish to consider candidate value functions that are functions of the finite event horizon, which is why it is not an exact step. In this approximate policy evaluation step, we define the value at an infinitely augmented state as a function of the state augmented by its finite event horizon, by sampling the previous older events from some distribution $\mu_{1}$. This $\mu_{1}$ is theoretically arbitrary and empirically can be based on the actual event process. The value function can thus be approximated using Monte Carlo methods. For the purpose of analysis, we consider the exact evaluation of the approximate function.
@@
Solving the optimization problem in the policy improvement step requires knowledge of the models $r(s,a,s^{\prime})$, $Q$ and $Q^{X}$. In this work, we ignore any approximation errors arising due to estimating the models and analyze the algorithm assuming the exact step is taken as defined.
@@ 6.1. Analysis
For the analysis of this algorithm, we need the following Lemma, which bounds the difference in the value of the value function at two different augmented states when they differ in events that are more than $T$ time steps old.
@@
$\left|V^{\pi}(\bar{s}_{1})-V^{\pi}(\bar{s}_{2})\right|\leq\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$ (1)
Throughout the rest of this section, we let $\epsilon=\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)$, the change in the value function induced by changing an arbitrary number of sufficiently old events.
Now, we characterize the behavior of the policy iteration procedure described previously. In general, exact policy iteration results in a sequence of policies that converge to the optimal policy. However, our approximate policy iteration produces policies whose value functions satisfy the following Lemma.
@@
That is, there is no guarantee that the sequence of value functions is non-decreasing. The only guarantee is that they do not decrease by more than the specific amount given by the Lemma. This is due to the approximation error $\epsilon$ at each time step.
@@
However, intuitively it would seem that such a small approximation error $\epsilon$ should only disturb the convergence of the algorithm when we are $\epsilon$-close to the optimal solution. During the initial iterations when we are very far from the solution, such small approximation errors should not matter.
It turns out that this is actually the case in some sense. Degradation of the value function occurs only in parts of the state space where the Bellman error is close to zero. A Bellman error of zero generally, although not always, corresponds to the optimal policy. This interplay between Bellman error and approximation error in our algorithm is formalized by the following result.
@@
At iteration $k$, for any augmented state $\bar{s}\in S\times\prod_{t=0}^{\infty}\mathcal{X}$, at least one of the following holds:
@@
    $|\mathcal{T}V^{\pi_{k}}(\bar{s})-V^{\pi_{k}}(\bar{s})|<\frac{11}{(1-\gamma)^{3}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$.
That is, the performance of the policy improves everywhere except the region of the state space where the Bellman error is very small. The faster the decay of exogenous event influence, the lesser the approximation error and smaller the region of state space where the policy may not improve.
@@
$<\max_{a\in A}\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}((s^{\prime},x^{\prime},x_{0:T-1}))\right]-\delta$
@@
Suppose the above condition holds. Then,
@@
$\qquad\qquad+\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+\gamma\epsilon+\gamma\epsilon+\epsilon-\delta$
@@
$\qquad\qquad+\frac{1}{1-\gamma}\left[\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+(1+2\gamma)\epsilon\right]$
@@
From this, it can be seen that as the time window $T$ considered increases, the error reduces as $e^{-\left\{\bar{\lambda},\lambda_{\alpha},\lambda_{\beta}\right\}T}$. So, depending on the rate of influence decay, increasing the time window beyond a certain point gives diminishing returns. The error also has a proportional dependence on $c$, the base intensity of the Hawkes process and an inverse dependency on $\lambda_{\alpha}$, the decay of the excitation function of the Hawkes process.
@@ 7. Sample Complexity
The algorithm given in Section 6 defines the value function and its update. But in practice, the value function has to be learnt using samples obtained from the environment. Further, the actual value function is generally approximated from a chosen set of functions.
In this section, we analyze the sample complexity of policy evaluation by using samples from a single sample path induced by the policy, and give error bounds on the evaluation error for the case of a linear function space and some regularity assumptions on the MDP.
@@
=$\mathcal{F}=\left\{f\;o\;f^{T}_{\text{trunc}}:f\in\mathcal{F}^{T}\right\},$
where $f_{\text{trunc}}:\bar{S}\rightarrow S\times\bigcup_{t=0}^{T}\mathcal{X}$ removes extra events that are older than $T$ time steps while preserving the current state and new events, and $\mathcal{F}^{T}$ is some function class of functions defined on this new truncated domain. For a fixed $f^{T}_{\text{trunc}}$, finding the best function in $\mathcal{F}$ is the same as finding the best function in $\mathcal{F}^{T}$.
@@
$\left|\left|V-\tilde{V}\right|\right|_{\rho},$
where $V$ is the true value function on $\bar{S}$, $\tilde{V}$ is the learned value function in $\mathcal{F}$, and $\rho$ is some distribution over $\bar{S}$, and the norm is the $l^{2}(\rho)$-norm, which is the expected $l^{2}$ norm of the value of the function w.r.t the measure $\rho$, i.e,
$\left|\left|f\right|\right|^{2}_{\rho}=\int_{\bar{S}}f(x)^{2}\;\rho(\mathrm{d}x).$
While the $f^{T}_{trunc}$ operator just ignores old events, we also define a projection operator $\Pi^{trunc}$ onto the function space defined on the truncated domain, as
$\Pi^{trunc}V=\inf_{f:S\times\bigcup_{t=0}^{T}\mathcal{X}}\left|\left|f-V\right|\right|_{\rho}.$
@@
A standard approximating function space considered is a an arbitrary finite dimensional function space. Specifically, assume that there are $d$ basis functions $\varphi_{i}:S\times\mathcal{X}^{T+1}\rightarrow\mathbb{R}$ that linearly span $\mathcal{F}$ and are bounded by $L$, that is, for each $f\in\mathcal{F}$, there exist $\alpha\in\mathbb{R}^{d}$ such that $f=\sum_{i\in[d]}\alpha_{i}\varphi_{i}$. These basis functions together form a feature representation function $\phi:S\times\mathcal{X}^{T+1}\rightarrow\mathbb{R}^{d}$, defined by $\phi(x)=(\varphi_{1}(x),\dots,\varphi_{d}(x))$.

### v3

@@ 3.1. Examples
Non-Markov self-exciting event processes that can be described as above are studied extensively in finance, economics, epidemiology, etc., in the form of variants of Hawkes and jump processes. While they are studied predominantly in continuous time and sometimes in discrete time, we consider discrete-time versions to fit into the traditional discrete-time reinforcement learning framework.
Consider a discrete-time marked Hawkes process $\left(E_{t},X_{t}\right)_{n\in\mathbb{N}}$ defined as follows: $\left(E_{t}\right)_{t\in\mathbb{N}}$ is a sequence of random variables taking values in $\left\{0,1\right\}$, with $E_{t}$ sampled from $\textsf{Bernoulli}\left(p_{t}\right)$, where the intensity $p_{t}$ at time $t$ depends on the realizations of $E_{t}$ at the previous time steps as:
@@
where $(\alpha_{t})$ and $(\beta_{t})$ are sequences that converge sufficiently quickly to zero. The sum $S_{t}=\sum_{t^{\prime}=1}^{t}E_{t}$ is a self-exciting counting process that is the discrete-time counterpart of the continuous-time Hawkes process.
The above-defined sequence of random variables satisfies Assumptions (A1) and (A2) required for the external process. $X_{t}$ is zero when no event occurs, and old events have a reduced influence on current events. More precisely, given two histories $(X_{t^{\prime}})_{t^{\prime}<t}$ and $(X^{\prime}_{t^{\prime}})_{t^{\prime}<t}$ that only differ by one event at some time $t^{\prime}\leq t-T$ with $E^{\prime}_{t^{\prime}}=X^{\prime}_{t^{\prime}}=0$ and $E_{t^{\prime}}=1$, $X_{t^{\prime}}=x$, the difference in intensities is $p_{t}-p^{\prime}_{t}=\alpha_{t-t^{\prime}}$, and hence,
@@
For sufficiently quickly converging $\alpha_{T}$ and $\beta_{T}$, the total variation perturbation $N_{T}$ due to events older than $T$ time steps goes to zero, satisfying Assumption (A2).
While $(\alpha_{n})$ and $(\beta_{n})$ are any general sequences that need to satisfy some regularity conditions (Seol, 2015) and the above bound, they could also decay in a more structured manner. For instance, for some $c,\lambda>0$ letting
@@
The details of notation that is used in the rest of the paper is provided in Table 1.
4. Guarantees for the existence of almost optimal solutions
Consider the MDP $\mathcal{M}=(S,A,r,\gamma,Q)$, where $S$ and $A$ are closed subsets of $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$ respectively, and $r:S\times A\times S\rightarrow\mathbb{R}$ is the reward as a function of the state, action, and resultant next state. Given an external discrete-time temporal process $\{(t_{i},X_{i})\}_{i}$ that acts on $\mathcal{M}$, we construct a new augmented MDP where the marks of all the events that occurred till time $t$ are appended to the state $s\in S$ to form a new augmented state $\bar{s}\in\bar{S}$, where
$\bar{s}=(s,x_{t},x_{t-1},x_{t-2},x_{t-3},\dots),\qquad\bar{S}=S\times\prod_{t=0}^{\infty}\mathcal{X}\left(\subseteq\mathbb{R}^{m}\times\prod_{t=0}^{\infty}\mathbb{R}^{d}\right).$
@@ 4. Guarantees for the existence of almost optimal solutions
Now, the state space of this MDP is infinite-dimensional and poses several challenges. However, per Assumptions A2 and A3 in Section 3, distant past events do not affect the transition probabilities during the current time step, with the total variation of the effect of events older than $t$ time steps on the transition dynamics and event distribution being bounded by $M_{t}$ and $N_{t}$ respectively. This allows one to study the possibility of finding a suitable policy that only depends on a finite horizon of past events. The following result formalizes this.
@@
Suppose the MDP and the external process satisfy the assumptions (A1-A3) and the regularity conditions (B1-B5). Then, we have the following.
@@
This result confirms that the formulated problem is well-posed, with an optimal solution that can be approximated using a tractable policy that is a function of the current state and only a finite history of events. This approximation can be as accurate as necessary based on the properties of the exogenous event process and the size of the event history considered.
The past event horizon $T$ above required for $\epsilon$-suboptimality is some $T$ that satisfies
@@
For each time $T>0$, we construct MDP $\mathcal{M}^{(T)}_{X}$ as having the same underlying stochastic process as $\mathcal{M}_{X}$, with only one difference: only events in the past $T$ time steps affect the current state transition, and all the events occurring before that have no effect and can effectively be replaced by zeroes. Depending on the context, the state space of this MDP can either be taken as $S\times\prod_{t=0}^{\infty}\mathcal{X}$ or $S\times\prod_{t=0}^{T}\mathcal{X}$, since all the older events are zeroed out.
Similarly, any policy $\pi$ on the extended state space $\bar{S}$ can operate in both $\mathcal{M}_{X}$ and $\mathcal{M}^{(T)}_{X}$ and thus induces different value functions $V^{\pi}_{\mathcal{M}_{X}}$ and $V^{\pi}_{\mathcal{M}^{(T)}_{X}}$ in these two MDPs respectively.
Since the MDP $\mathcal{M}_{X}$ is the actual MDP under consideration and $\mathcal{M}^{(T)}_{X}$ is just an approximation, when there is no ambiguity, we use just $V^{\pi}$ without any subscript to denote the value function of policy $\pi$ in the augmented non-stationary MDP $\mathcal{M}_{X}$ and use subscripts when required to distinguish between multiple MDPs.
@@
Let $\pi$ be an arbitrary policy as a function of the augmented state $\bar{s}\in\bar{S}$. That is, $\pi$ can be either stochastic or deterministic and can depend on the current state $s\in S$ and any number of past events in $\mathcal{X}$. Then,
$\left|\left|V^{\pi}_{\mathcal{M}^{(T)}_{X}}-V^{\pi}_{\mathcal{M}_{X}}\right|\right|_{\infty}\leq\frac{1}{1-\gamma}\left(\left\|r\right\|_{\infty}+\gamma\left\|V^{\pi}_{\mathcal{M}^{(T)}_{X}}\right\|_{\infty}\right)\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right),$
where $r$ is the reward function and $V^{\pi}_{\mathcal{M}}$ denotes the value function of policy $\pi$ in MDP $\mathcal{M}$.
Proof.
See Appendix B.1. ∎
@@
Part (1) follows from the assumptions and Theorem 4.2.3 in Hernández-Lerma and Lasserre (1996). For part (2), let $\pi^{*}$ be the deterministic optimal stationary policy for $\mathcal{M}_{X}$ and let $\pi^{*(T)}$ be the deterministic optimal stationary policy for $\mathcal{M}^{(T)}$. From Lemma 4.2, since $\|r\|_{\infty}\leq 1$ and $\|V\|_{\infty}\leq\frac{1}{1-\gamma}$,
$V^{\pi^{*(T)}}_{\mathcal{M}_{X}}$ $\geq V^{\pi^{*(T)}}_{\mathcal{M}^{(T)}_{X}}-\frac{1}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
$\geq V^{\pi^{*}}_{\mathcal{M}^{(T)}_{X}}-\frac{1}{(1-\gamma)^{2}}\left(\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right)\right)$
@@
Choosing such a $T$ gives the Theorem. ∎
@@ 5. A policy iteration Algorithm
Now that an $\epsilon$-optimal policy exists that is a function of only a finite event horizon, we wish to find such a policy. We propose a policy iteration algorithm that alternates between approximate policy evaluation and approximate policy improvement. This procedure is listed in Algorithm 1.
@@
For the policy evaluation step, we only wish to consider candidate value functions that are functions of the finite event horizon, which is why it is not an exact step. We define the value function in $\mathcal{M}_{X}$ at an infinitely augmented state as a function of the state augmented by just a finite event horizon by sampling the previous older events from some arbitrary distribution $\mu_{1}$. In practice, $\mu_{1}$ is just the actual event process. The value function can thus be approximated using Monte Carlo methods. For the purpose of analysis, we consider the exact evaluation of this approximate value function.
@@
Solving the optimization problem in the policy improvement step requires knowledge of the models $r(s,a,s^{\prime})$, $Q$, and $Q^{X}$. In this work, we ignore any approximation errors arising due to estimating the models and analyze the algorithm by assuming the exact step is taken as defined.
@@ 5.1. Analysis
For the analysis of this algorithm, we need the following Lemma, which bounds the difference in the value function at two different augmented states when they differ in events that are more than $T$ time steps old.
@@
Proof.
See Appendix B.2. ∎
Here, $(s,x_{0:T},(0)_{T+1:\infty})=(s,x_{0},x_{1},\dots,x_{T},0,0,\dots)\in\bar{S}$ is the extended state that has been essentially “truncated” and made finite by replacing events older than $T$ time steps by zero.
@@
$\left|V^{\pi}(\bar{s}_{1})-V^{\pi}(\bar{s}_{2})\right|\leq\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}\left(M_{t}+N_{t}\right).$
Essentially, this means that if a policy considers only events in the past $T$ time steps, its value at states differing at older events differ by an amount proportional to the extent of non-stationarity induced by all events older than $T$ steps.
This helps characterize the behavior of the policy iteration procedure described previously. While exact policy iteration results in a sequence of policies that converge to the optimal policy, our approximate policy iteration produces policies whose value functions satisfy the following Lemma, whose proof is given later in this section.
@@
That is, while there is no guarantee that the sequence of value functions $\left\{V^{\pi_{k}}\right\}$ is non-decreasing, Lemma 5.3 provides a guarantee that they do not decrease by more than a specific amount. This is due to the approximation error
$\epsilon=\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$
at each time step induced by the non-stationarity due to the exogenous events.
@@
Intuitively it would seem that such a small approximation error $\epsilon$ should only disturb the convergence of the algorithm when we are $\epsilon$-close to the optimal solution. During the initial iterations, when we are very far from the solution, such small approximation errors should not matter.
We next establish theoretically that the degradation of the value function occurs only in parts of the state space where the Bellman error is close to zero. A Bellman error of zero generally, although not always, corresponds to the optimal policy. This interplay between Bellman error and approximation error in our algorithm is formalized by the following result.
@@
At iteration $k$ of the policy iteration procedure given in Algorithm 1, for any augmented state $\bar{s}\in\bar{S}=S\times\prod_{t=0}^{\infty}\mathcal{X}$, at least one of the following holds:
@@
    $|\mathcal{T}V^{\pi_{k}}(\bar{s})-V^{\pi_{k}}(\bar{s})|<\frac{11}{(1-\gamma)^{3}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$,
where $V^{\pi}$ denotes the value function of policy $\pi$ in the MDP $\mathcal{M}_{X}$.
That is, the performance of the policy improves everywhere except the region of the state space where the Bellman error is very small. The faster the decay of exogenous event influence, the lesser the approximation error and the smaller the region of state space where the policy may not improve.
@@
$\qquad\qquad\qquad<\max_{a\in A}\mathsf{E}_{\begin{subarray}{c}s^{\prime}\sim Q_{x_{0:T},x^{\prime}_{T+1:\infty}}(.|s,a)\\ x^{\prime}\sim Q^{X}_{x_{0:T},x^{\prime}_{T+1:\infty}}(.)\\ x^{\prime}_{T+1:\infty}\sim\mu_{2}\end{subarray}}\left[r(s,a,s^{\prime})+\gamma\hat{V}_{k}((s^{\prime},x^{\prime},x_{0:T-1}))\right]-\delta$
@@
Suppose the above condition holds. Then, for $\epsilon=\frac{2}{(1-\gamma)^{2}}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})$,
@@
$\qquad\qquad+\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+\gamma\epsilon+\gamma\epsilon+\epsilon-\delta.$
@@
$\qquad\qquad\qquad\qquad+\frac{1}{1-\gamma}\left[\frac{1}{1-\gamma}\sum_{t=T+1}^{\infty}(M_{t}+N_{t})+(1+2\gamma)\epsilon\right]$
@@
From this, it can be seen that as the time window $T$ considered increases, the error reduces as $e^{-\left\{\bar{\lambda},\lambda_{\alpha},\lambda_{\beta}\right\}T}$. Hence, depending on the rate of influence decay, increasing the time window beyond a certain point gives diminishing returns. The error also has a proportional dependence on $c$, the base intensity of the Hawkes process and an inverse dependency on $\lambda_{\alpha}$, the decay of the excitation function of the Hawkes process.
@@ 6. Sample Complexity
While the algorithm given in Section 5 defines the value function and its update, in practice, it has to be learnt using samples obtained from the environment and is generally approximated from a chosen class of functions. In this section, we analyze the sample complexity of policy evaluation by pathwise Least-Squares Temporal Difference (LSTD) (Lazaric et al., 2012), which uses samples from a single sample path induced by the policy, and give error bounds on the evaluation error for the case of a linear function space and some regularity assumptions on the MDP.
In contrast to Lazaric et al. (2012), our work considers an MDP with an infinite-dimensional augmented state space, a stochastic reward function and an additional error due to the use of tractable policies and value functions on a finite-dimensional domain. This results in additional sample complexity even for evaluating a given policy, depending on the extent to which the exogenous events affect the evolution of the states. In this section, we quantify the precise nature of this error.
@@
=$\mathcal{F}=\left\{f\;o\;f^{(T)}_{\text{trunc}}:f\in\mathcal{F}^{(T)}\right\},$
where $f^{(T)}_{\text{trunc}}:\bar{S}\rightarrow S\times\bigcup_{t=0}^{T}\mathcal{X}$ removes extra events that are older than $T$ time steps while preserving the current state and new events, and $\mathcal{F}^{(T)}$ is some function class of functions defined on this new truncated domain. For a fixed $T$, finding the best function in $\mathcal{F}$ is the same as finding the best function in $\mathcal{F}^{(T)}$.
@@
$\left\|V-\widetilde{V}\right\|_{\rho},$
where $V$ is the true value function on $\bar{S}$, $\widetilde{V}$ is the learned value function in $\mathcal{F}$, and $\rho$ is some distribution over $\bar{S}$, and the norm is the $l^{2}(\rho)$-norm, which is the expected $l^{2}$ norm of the value of the function w.r.t the measure $\rho$, i.e,
$\left\|f\right\|^{2}_{\rho}=\int_{\bar{S}}f(x)^{2}\;\rho(\mathrm{d}x).$
While the $f^{(T)}_{trunc}$ operator just ignores old events, we also define a projection operator $\Pi^{trunc}$ onto the function space defined on the truncated domain, as
$\Pi^{trunc}V=\inf_{f:S\times\bigcup_{t=0}^{T}\mathcal{X}\rightarrow\mathbb{R}}\left\|f-V\right\|_{\rho}.$
@@
A standard approximating function space considered is a an arbitrary finite dimensional function space. Specifically, assume that there are $d$ basis functions $\varphi_{i}:S\times\mathcal{X}^{T+1}\rightarrow\mathbb{R}$ that linearly span $\mathcal{F}^{(T)}$ and are bounded by $L$, that is, for each $f\in\mathcal{F}^{(T)}$, there exist $\alpha\in\mathbb{R}^{d}$ such that $f=\sum_{i\in[d]}\alpha_{i}\varphi_{i}$. These basis functions together form a feature representation function $\phi:S\times\mathcal{X}^{T+1}\rightarrow\mathbb{R}^{d}$, defined by $\phi(x)=(\varphi_{1}(x),\dots,\varphi_{d}(x))$.